OpenStudy (anonymous):

Find the domain of each function. (a) y = log4(7x) (b) y = log10(3 − 8x) (c) y = ln(x2) (d) y = (ln(x))2 (e) y = ln(x2 − 49)

OpenStudy (anonymous):

c. ln(x^2) d. y=ln((x))^2

OpenStudy (anonymous):

rembr x cant be 0 or 1as its log

OpenStudy (anonymous):

For all logarithms, the domain is anywhere the argument is <0. So just set the arguments all to zero. For example, the first one would be x>0. The second would be 3-8x>0 3>8x x<3/8

OpenStudy (anonymous):

thanks that is easy to remember

OpenStudy (anonymous):

Anywhere the the argument is >0. That is a mistype on my part. And logs can be 1 Ishaan. For example e^0=1 So ln(1)=ln(e^0)=0.